# Global Optimal Closed-Form Solutions for Intelligent Surfaces With Mutual Coupling: Is Mutual Coupling Detrimental or Beneficial?

### v1

@@ Abstract
Index Terms:
Beyond diagonal RIS (BD-RIS), multiport network theory, mutual coupling, reconfigurable intelligent surface (RIS).
@@ I Introduction
Despite recent efforts in studying RIS with mutual coupling, two open challenges can be identified. First, existing mutual coupling-aware optimization algorithms suffer from a high computational complexity caused by the high number of iterations required to converge, which becomes prohibitive when the number of RIS elements increases. Second, given their iterative nature, their convergence is guaranteed only to a local optimum. Thus, an expression of the achievable channel gain of a RIS-aided system with mutual coupling remains unknown. To solve these two challenges, in this study, we derive global optimal closed-form solutions to maximize the channel gain of BD-RIS-aided systems with mutual coupling, applicable to the fully- and tree-connected RIS architectures. In addition, we characterize and compare the scaling laws of the channel gain in the presence and in the absence of mutual coupling. The contributions of this study can be summarized as follows.
@@
Notation: Vectors and matrices are denoted with bold lower and bold upper letters, respectively. Scalars are represented with letters not in bold font. $\Re\{a\}$, $\Im\{a\}$, $|a|$, $\arg(a)$, and $a^{*}$ refer to the real part, imaginary part, modulus, phase, and conjugate of a complex scalar $a$, respectively. $[\mathbf{a}]_{i}$ and $\|\mathbf{a}\|_{2}$ refer to the $i$th element and $l_{2}$-norm of a vector $\mathbf{a}$, respectively. $\mathbf{A}^{*}$, $\mathbf{A}^{T}$, $\mathbf{A}^{H}$, and $[\mathbf{A}]_{i,j}$ refer to the conjugate, transpose, conjugate transpose, and $(i,j)$th element of a matrix $\mathbf{A}$, respectively. $\mathbb{R}$ and $\mathbb{C}$ denote the real and complex number sets, respectively. $j=\sqrt{-1}$ denotes the imaginary unit. $\mathbf{0}$ and $\mathbf{I}$ denote an all-zero matrix and an identity matrix, respectively, with appropriate dimensions. $\mathcal{CN}(0,\sigma^{2})$ denotes the distribution of a circularly-symmetric complex Gaussian (CSCG) random variable whose real and imaginary parts are independent and Gaussian distributed with mean zero and variance $\sigma^{2}/2$. $\mathcal{CN}(\mathbf{0},\mathbf{C})$ denotes the distribution of a CSCG random vector with mean vector $\mathbf{0}$ and covariance matrix $\mathbf{C}$. diag$(a_{1},\ldots,a_{N})$ refers to a diagonal matrix with diagonal elements being $a_{1},\ldots,a_{N}$.
@@ II RIS-Aided Channel Model With
Consider a communication system between a single-antenna transmitter and a single-antenna receiver, aided by an $N_{I}$-element RIS. This system can be modeled by using the multiport network theory [3, 12, 14], and its wireless channel can be regarded as an $N$-port network, with $N=2+N_{I}$, as represented in Fig. 1.
@@
To obtain a tractable expression of the channel $h\in\mathbb{C}$ relating the voltage $v_{T}$ (transmitted signal) and the voltage $v_{R}$ (received signal) through $v_{R}=hv_{T}$, we make the following two assumptions, commonly considered in related literature [3, 12, 14]. First, the transmission distances from the transmitter to RIS, from the RIS to receiver, and from the transmitter to receiver are assumed to be large enough such that we can neglect the effect of the feedback channels, i.e., we can consider $\mathbf{z}_{TI}=\mathbf{0}$, $\mathbf{z}_{IR}=\mathbf{0}$, and $z_{TR}=0$, which is also known as the unilateral approximation [31]. Second, the antennas at the transmitter and receiver are assumed to be perfectly matched to $Z_{0}$, i.e., $z_{TT}=Z_{0}$ and $z_{RR}=Z_{0}$. With these two assumptions, it is possible to show that the channel $h$ based on the $Z$-parameters representation reads as
@@
where $Y_{0}=Z_{0}^{-1}$, $\mathbf{Y}_{I}=\mathbf{Z}_{I}^{-1}$, $\mathbf{Y}_{II}=\mathbf{Z}_{II}^{-1}$. Although (6) is equal to (2) given the equivalence between $Z$- and $Y$-parameters, the two models are best suited for their respective use cases. Specifically, the $Z$-parameters model in (2) is suitable for fully-connected RISs with a full matrix $\mathbf{Z}_{I}$. Meanwhile, (6) is suitable for tree-connected RISs, where the admittance matrix $\mathbf{Y}_{I}$ can explicitly capture the specific circuit topology of such RIS architectures. In the following sections, we derive global optimal solutions to maximize the channel gain for fully- and tree-connected RISs, using the $Z$- and $Y$-parameters representations, respectively.
@@ III-A Problem Formulation and Global Optimal Solution
For a lossless fully-connected RIS, $\mathbf{Z}_{I}$ is a purely imaginary matrix, i.e., $\mathbf{Z}_{I}=j\mathbf{X}_{I}$, where $\mathbf{X}_{I}\in\mathbb{R}^{N_{I}\times N_{I}}$ is the reactance matrix of the RIS. Besides, the reciprocity constraint imposes that $\mathbf{X}_{I}$ is symmetric, i.e., $\mathbf{X}_{I}=\mathbf{X}_{I}^{T}$. Given these constraints, the channel gain $|h|^{2}$ maximization problem is given by
@@
by using the channel model in (2).
@@
which is a real matrix since $\mathbf{X}_{I}$, $\Im\{\mathbf{Z}_{II}\}$, and $\Re\{\mathbf{Z}_{II}\}^{-1/2}$ are real matrices. Note that $\Re\{\mathbf{Z}_{II}\}^{-1/2}$ is a real matrix since $\Re\{\mathbf{Z}_{II}\}$ is positive semi-definite following Proposition 1 and its invertibility. Thus, by substituting
@@
×$\left.\times\left(j\bar{\mathbf{X}}_{I}+Z_{0}\mathbf{I}\right)^{-1}\sqrt{Z_{0}}\Re\{\mathbf{Z}_{II}\}^{-1/2}\mathbf{z}_{IT}\right|^{2}$ (11)
$\mathsf{\mathrm{s.t.}}\;\;\;$ (9)$,\;\mathbf{X}_{I}=\mathbf{X}_{I}^{T}.$ (12)
@@
To solve problem (13)-(14) with the solution proposed in [10], we map the $Z$-parameter based channels into the $S$-parameter based ones [14] by defining
@@
$\mathsf{\mathrm{s.t.}}\;\;\;$ (19)$,\;\bar{\mathbf{X}}_{I}=\bar{\mathbf{X}}_{I}^{T}.$ (21)
Noticing that constraint (21) implies that $\bar{\boldsymbol{\Theta}}$ can be an arbitrary unitary and symmetric matrix, we can equivalently transform (20)-(21) into
@@
Third, the global optimal $\mathbf{X}_{I}$ is obtained from $\bar{\mathbf{X}}_{I}^{\star}$ via (10).
@@ III-B Channel Gain Upper Bound
being an explicit function of the channels and the mutual coupling. In (27), the term $|z_{RT}-\mathbf{z}_{RI}\Re\{\mathbf{Z}_{II}\}^{-1}\mathbf{z}_{IT}/2|$ represents the strength of the direct link $z_{RT}$ and the structural scattering (or specular reflection) of the RIS $-\mathbf{z}_{RI}\Re\{\mathbf{Z}_{II}\}^{-1}\mathbf{z}_{IT}/2$. In addition, $\|\mathbf{z}_{RI}\Re\{\mathbf{Z}_{II}\}^{-1/2}\|_{2}$ and $\|\Re\{\mathbf{Z}_{II}\}^{-1/2}\mathbf{z}_{IT}\|_{2}$ are the strengths of the effective RIS-receiver and transmitter-RIS channels, respectively. Note that (27) is an explicit expression of the maximum channel gain achievable by a fully-connected RIS with mutual coupling, which is crucial to conducting performance analysis of RIS-aided systems with mutual coupling.
The fully-connected RIS is the most flexible BD-RIS architecture, enabling the highest performance at the cost of a higher circuit complexity. Specifically, in a fully-connected RIS, there are $N_{I}(N_{I}+1)/2$ tunable impedance components interconnecting all the RIS elements to each other and to ground [3]. To decrease the circuit complexity, the tree-connected RIS has been proposed, which is the least complex BD-RIS architecture achieving the same performance as the fully-connected RIS without mutual coupling, while including only $12N_{I}-$ 1 tunable impedance components [4]. In the following, we show that tree-connected RIS can achieve the same performance as the fully-connected RIS also in the presence of mutual coupling.
@@ IV-A Problem Formulation and Global Optimal Solution
For a lossless tree-connected RIS, its admittance matrix $\mathbf{Y}_{I}$ is purely imaginary, i.e., $\mathbf{Y}_{I}=j\mathbf{B}_{I}$, where $\mathbf{B}_{I}\in\mathbb{R}^{N_{I}\times N_{I}}$ is the susceptance matrix of the RIS. For a reciprocal RIS, $\mathbf{B}_{I}$ is symmetric, i.e., $\mathbf{B}_{I}=\mathbf{B}_{I}^{T}$. In addition, $\mathbf{B}_{I}$ is also subject to the constraints given by the specific tree-connected RIS architecture considered. Since numerous tree-connected RIS architectures are possible, we consider the tridiagonal RIS architecture in the following, having $[\mathbf{B}_{I}]_{i,j}=\mathbf{0}$ if $\left|i-j\right|>1$ [4]. Nevertheless, the following discussion is valid for every tree-connected RIS architecture. For a tridiagonal RIS, the channel gain $|h|^{2}$ maximization problem writes as
@@
$\mathsf{\mathrm{s.t.}}\;\;\;$ (30)$,\;\mathbf{B}_{I}=\mathbf{B}_{I}^{T},\;[\mathbf{B}_{I}]_{i,j}=\mathbf{0}\text{ if }\left|i-j\right|>1,$ (33)
where we introduced
@@
$\mathsf{\mathrm{s.t.}}\;\;\;$ (38),(30)$,\;\mathbf{B}_{I}=\mathbf{B}_{I}^{T},\;[\mathbf{B}_{I}]_{i,j}=\mathbf{0}\text{ if }\left|i-j\right|>1.$ (40)
@@
$\boldsymbol{\alpha}$ =$=\Re\{\mathbf{Y}_{II}\}^{-1/2}\bar{\boldsymbol{\alpha}},$ (47)
@@
Note that condition (46) is a system of $N_{I}$ linear equations in complex coefficients, i.e., the entries of $\boldsymbol{\alpha}$ and $\boldsymbol{\beta}$, with $12N_{I}-$ 1 real unknown, i.e., the entries of $\mathbf{B}_{I}$ that are not constrained to zero. Remarkably, it is possible to prove that the system in (46) has exactly one solution in general [4], and can be solved through the algorithm proposed in [4] for given $\boldsymbol{\alpha}$ and $\boldsymbol{\beta}$.
@@ V-A Scaling Law with Mutual Coupling
Third, to compute $\mathbb{E}[|\mathbf{z}_{RI}\Re\{\mathbf{Z}_{II}\}^{-1}\mathbf{z}_{IT}|]$ in (52), we use the eigenvalue decomposition of $\Re\{\mathbf{Z}_{II}\}^{-1}$. Since $\Re\{\mathbf{Z}_{II}\}^{-1}$ is a real symmetric matrix, we have $\Re\{\mathbf{Z}_{II}\}^{-1}=\mathbf{Q}\mathbf{\Lambda}\mathbf{Q}^{T}$, where $\mathbf{Q}\in\mathbb{R}^{N_{I}\times N_{I}}$ is an orthogonal matrix whose columns are the eigenvectors of $\Re\{\mathbf{Z}_{II}\}^{-1}$, and $\mathbf{\Lambda}=\text{diag}(\lambda_{1},\ldots,\lambda_{N_{I}})$, with $\lambda_{n_{I}}\in\mathbb{R}$ being the $n_{I}$th eigenvalue of $\Re\{\mathbf{Z}_{II}\}^{-1}$. Consequently, the random variable $R=\mathbf{z}_{RI}\Re\{\mathbf{Z}_{II}\}^{-1}\mathbf{z}_{IT}$ can be rewritten as $R=\mathbf{q}_{RI}\mathbf{\Lambda}\mathbf{q}_{IT}$, where $\mathbf{q}_{RI}=\mathbf{z}_{RI}\mathbf{Q}$ is distributed as $\mathbf{q}_{RI}\sim\mathcal{CN}(\mathbf{0},\rho_{RI}\mathbf{I})$ and $\mathbf{q}_{IT}=\mathbf{Q}^{T}\mathbf{z}_{IT}$ is distributed as $\mathbf{q}_{IT}\sim\mathcal{CN}(\mathbf{0},\rho_{IT}\mathbf{I})$ due to the orthogonality of $\mathbf{Q}$. Since $R$ can be expressed as $R=\mathbf{q}_{RI}\mathbf{\Lambda}\mathbf{q}_{IT}$, $R$ is the sum of $N_{I}$ independent random variables, i.e., $R=\sum_{n_{I}=1}^{N_{I}}R_{n_{I}}$, where $R_{n_{I}}=[\mathbf{q}_{RI}]_{n_{I}}\lambda_{n_{I}}[\mathbf{q}_{IT}]_{n_{I}}$ has variance $\rho_{RI}\rho_{IT}\lambda_{n_{I}}^{2}$, for $n_{I}=1,\ldots,N_{I}$. Thus, according to the Lyapunov central limit theorem (CLT), $R$ is distributed as $R\sim\mathcal{CN}(0,\rho_{RI}\rho_{IT}\sum_{n_{I}=1}^{N_{I}}\lambda_{n_{I}}^{2})$. Since $\lambda_{n_{I}}^{2}$, for $n_{I}=1,\ldots,N_{I}$, are the eigenvalues of $\Re\{\mathbf{Z}_{II}\}^{-2}$, we have $\sum_{n_{I}=1}^{N_{I}}\lambda_{n_{I}}^{2}=\text{Tr}(\Re\{\mathbf{Z}_{II}\}^{-2})$, giving $R\sim\mathcal{CN}(0,\rho_{RI}\rho_{IT}\text{Tr}\left(\Re\{\mathbf{Z}_{II}\}^{-2}\right))$ and

### v2

@@ I Introduction
Despite recent efforts in studying RIS with mutual coupling, two open challenges can be identified. First, existing mutual coupling-aware optimization algorithms suffer from a high computational complexity caused by the high number of iterations required to converge, which becomes prohibitive when the number of RIS elements increases. Second, given their iterative nature, their convergence is guaranteed only to a local optimum. Thus, an expression of the achievable channel gain of a RIS-aided system with mutual coupling remains unknown. To solve these two challenges, in this study, we derive global optimal closed-form solutions to maximize the channel gain of BD-RIS-aided systems with mutual coupling, applicable to the fully- and tree-connected RIS architectures. In addition, we characterize and compare the scaling laws of the channel gain in the presence and in the absence of mutual coupling. Different from conventional efforts relying on iterative solutions converging to locally optimal solutions, we adopt a different strategy in this paper. We first reformulate the objective function and the constraints in an equivalent way by introducing auxiliary variables. We then derive an upper bound on the reformulated objective function and provide a closed-form solution for the optimization variables able to exactly achieve the upper bound. The contributions of this study can be summarized as follows.
@@
[…]
@@ II RIS-Aided Channel Model With
Consider a communication system between a single-antenna transmitter and a single-antenna receiver, aided by an $N_{I}$-element RIS. This system can be modeled by using the multiport network theory [3, 12, 14], and its wireless channel can be regarded as an $N$-port network, with $N=2+N_{I}$, as represented in Fig. 1. We consider a single-input single-output (SISO) system to gain fundamental insights into the role of mutual coupling in RIS-aided wireless systems, as in related literature [15, 16, 29]. Nevertheless, the illustrated channel model can be readily extended to multiple-input multiple-output (MIMO) systems [14].
@@
To obtain a tractable expression of the channel $h\in\mathbb{C}$ relating the voltage $v_{T}$ (transmitted signal) and the voltage $v_{R}$ (received signal) through
$v_{R}=hv_{T}+n,$ (2)
where $n$ is the additive white Gaussian noise (AWGN), we make the following two assumptions, commonly considered in related literature [3, 12, 14]. First, the transmission distances from the transmitter to RIS, from the RIS to receiver, and from the transmitter to receiver are assumed to be large enough such that we can neglect the effect of the feedback channels, i.e., we can consider $\mathbf{z}_{TI}=\mathbf{0}$, $\mathbf{z}_{IR}=\mathbf{0}$, and $z_{TR}=0$, which is also known as the unilateral approximation [31]. Second, the antennas at the transmitter and receiver are assumed to be perfectly matched to $Z_{0}$, i.e., $z_{TT}=Z_{0}$ and $z_{RR}=Z_{0}$. With these two assumptions, it is possible to show that the channel $h$ based on the $Z$-parameters representation reads as
@@
where $Y_{0}=Z_{0}^{-1}$, $\mathbf{Y}_{I}=\mathbf{Z}_{I}^{-1}$, $\mathbf{Y}_{II}=\mathbf{Z}_{II}^{-1}$. Although (7) is equal to (3) given the equivalence between $Z$- and $Y$-parameters (as shown in [14]), the two models are best suited for their respective use cases. Specifically, the $Z$-parameters model in (3) is suitable for fully-connected RISs with a full matrix $\mathbf{Z}_{I}$. For the sake of illustration, we report the diagram of a $4$-element fully-connected RIS in Fig. 2(a), showing that each antenna port is connected to ground and to all other ports through a tunable impedance [3]. Meanwhile, (7) is suitable for tree-connected RISs, where the admittance matrix $\mathbf{Y}_{I}$ can explicitly capture the specific circuit topology of such RIS architectures. The diagram of a $4$-element tree-connected RIS is shown in Fig. 2(b), where the antenna ports are interconnected to each other such that the resulting topology can be represented with an acyclic and connected graph, i.e., a tree [4]. In the following sections, we derive global optimal solutions to maximize the channel gain for fully- and tree-connected RISs, using the $Z$- and $Y$-parameters representations, respectively. The channel gain maximization problem is considered since it is equivalent to maximizing the signal-to-noise ratio (SNR) at the receiver.
@@ III Optimizing Fully-Connected RIS With
Figure 2: Circuit topology of a (a) fully-connected RIS and (b) tree-connected RIS, with $N_{I}=4$ elements.
@@ III-A Problem Formulation and Global Optimal Solution
For a lossless fully-connected RIS, $\mathbf{Z}_{I}$ is a purely imaginary matrix, i.e., $\mathbf{Z}_{I}=j\mathbf{X}_{I}$, where $\mathbf{X}_{I}\in\mathbb{R}^{N_{I}\times N_{I}}$ is the reactance matrix of the RIS. Besides, the reciprocity of the microwave network implementing the fully-connected RIS imposes that $\mathbf{X}_{I}$ is symmetric, i.e., $\mathbf{X}_{I}=\mathbf{X}_{I}^{T}$. Given these constraints, the channel gain $|h|^{2}$ maximization problem is given by
@@
by using the channel model in (3). To solve (8)-(9), we assume that the mutual coupling matrix $\mathbf{Z}_{II}$ is known. This matrix depends purely on the array geometry, i.e., antenna element type and inter-element distances, and is independent of the channels between the RIS and the transmitter and receiver. Thus, it can be estimated offline, before the RIS is deployed. We also assume perfect CSI, which can be acquired using the channel estimation protocols developed in previous literature [21, 22], suitably adapted for BD-RIS.
@@
which is a real matrix since $\mathbf{X}_{I}$, $\Im\{\mathbf{Z}_{II}\}$, and $\Re\{\mathbf{Z}_{II}\}^{-1/2}$ are real matrices. Note that $\Re\{\mathbf{Z}_{II}\}^{-1/2}$ is a real matrix since $\Re\{\mathbf{Z}_{II}\}$ is positive semi-definite following Proposition 1 and its invertibility. The rationale behind this auxiliary variable is that it allows to “diagonalize” the mutual coupling, similar to the effect of the decoupling network used in [24]. Thus, by substituting
@@
$\left.\times\left(j\bar{\mathbf{X}}_{I}+Z_{0}\mathbf{I}\right)^{-1}\sqrt{Z_{0}}\Re\{\mathbf{Z}_{II}\}^{-1/2}\mathbf{z}_{IT}\right|^{2}$ (12)
[…]
@@
Interestingly, (14) is equivalent to the channel expression in the presence of a RIS with reactance matrix $\bar{\mathbf{X}}_{I}$ and no mutual coupling, and a similar expression has been obtained through the use of a decoupling network in [24]. To solve problem (14)-(15) with the solution proposed in [10], we map the $Z$-parameter based channels into the $S$-parameter based ones [14] by defining
@@
[…]
Noticing that constraint (22) implies that $\bar{\boldsymbol{\Theta}}$ can be an arbitrary unitary and symmetric matrix (since $\bar{\mathbf{X}}_{I}$ is a real matrix given the positive-definiteness of $\Re\{\mathbf{Z}_{II}\}$), we can equivalently transform (21)-(22) into
@@
Third, the global optimal $\mathbf{X}_{I}$ is obtained from $\bar{\mathbf{X}}_{I}^{\star}$ via (11). The computational complexity of optimizing a fully-connected RIS is driven by the complexity of solving problem (23)-(24) with the solution proposed in [10]. Since the solution in [10] requires the computation of the eigenvalue decomposition of an $N_{I}\times N_{I}$ matrix, its complexity scales with $\mathcal{O}(N_{I}^{3})$.
@@ III-B Channel Gain Upper Bound
being an explicit function of the channels and the mutual coupling. In (28), the term $|z_{RT}-\mathbf{z}_{RI}\Re\{\mathbf{Z}_{II}\}^{-1}\mathbf{z}_{IT}/2|$ represents the strength of the direct link $z_{RT}$ and the structural scattering (or specular reflection) of the RIS $\mathbf{z}_{RI}\Re\{\mathbf{Z}_{II}\}^{-1}\mathbf{z}_{IT}/2$. In addition, $\|\mathbf{z}_{RI}\Re\{\mathbf{Z}_{II}\}^{-1/2}\|_{2}$ and $\|\Re\{\mathbf{Z}_{II}\}^{-1/2}\mathbf{z}_{IT}\|_{2}$ are the strengths of the effective RIS-receiver and transmitter-RIS channels, respectively. Note that (28) is an explicit expression of the maximum channel gain achievable by a fully-connected RIS with mutual coupling, which is crucial to conducting performance analysis of RIS-aided systems with mutual coupling.
The fully-connected RIS is the most flexible BD-RIS architecture, enabling the highest performance at the cost of a higher circuit complexity. Specifically, in a fully-connected RIS, there are $N_{I}(N_{I}+1)/2$ tunable impedance components interconnecting all the RIS elements to each other and to ground [3]. To decrease the circuit complexity, the tree-connected RIS has been proposed, which is the least complex BD-RIS architecture achieving the same performance as the fully-connected RIS without mutual coupling, while including only $2N_{I}-1$ tunable impedance components [4]. In the following, we show that tree-connected RIS can achieve the same performance as the fully-connected RIS also in the presence of mutual coupling.
@@ IV-A Problem Formulation and Global Optimal Solution
For a lossless tree-connected RIS, its admittance matrix $\mathbf{Y}_{I}$ is purely imaginary, i.e., $\mathbf{Y}_{I}=j\mathbf{B}_{I}$, where $\mathbf{B}_{I}\in\mathbb{R}^{N_{I}\times N_{I}}$ is the susceptance matrix of the RIS. For a reciprocal RIS, $\mathbf{B}_{I}$ is symmetric, i.e., $\mathbf{B}_{I}=\mathbf{B}_{I}^{T}$. In addition, $\mathbf{B}_{I}$ is also subject to the constraints given by the specific tree-connected RIS architecture considered. Since numerous tree-connected RIS architectures are possible, we consider the tridiagonal RIS architecture in the following, having $\mathbf{B}_{I}]_{i,j}=\mathbf{0}$ if $\left|i-j\right|>1$ [4]. Nevertheless, the following discussion is valid for every tree-connected RIS architecture. For a tridiagonal RIS, the channel gain $|h|^{2}$ maximization problem writes as
@@
$\underset{\mathbf{B}_{I}}{\mathsf{\mathrm{max}}}\;\;$ $\frac{1}{4Y_{0}^{2}}\left|y_{RT}-\mathbf{y}_{RI}\Re\{\mathbf{Y}_{II}\}^{-1/2}\sqrt{Y_{0}}\right$.
$\left.\times\left(j\bar{\mathbf{B}}_{I}+Y_{0}\mathbf{I}\right)^{-1}\sqrt{Y_{0}}\Re\{\mathbf{Y}_{II}\}^{-1/2}\mathbf{y}_{IT}\right|^{2}$ (33)
[…]
which can be rewritten as
@@
[…]
by introducing
@@
[…]
@@
$\boldsymbol{\alpha}$ $=\Re\{\mathbf{Y}_{II}\}^{-1/2}\bar{\boldsymbol{\alpha}},$ (50)
@@
Note that condition (49) is a system of $N_{I}$ linear equations in complex coefficients, i.e., the entries of $\boldsymbol{\alpha}$ and $\boldsymbol{\beta}$, with $2N_{I}-1$ real unknown, i.e., the entries of $\mathbf{B}_{I}$ that are not constrained to zero. Remarkably, it is possible to prove that the system in (49) has exactly one solution in general [4], and can be solved through the algorithm proposed in [4] for given $\boldsymbol{\alpha}$ and $\boldsymbol{\beta}$. The computational complexity of optimizing a tree-connected RIS is driven by the complexity of solving the system of equations in (49) with the solution proposed in [4]. Since the system of equations in (49) includes $2N_{I}-1$ independent equations with real coefficients and $2N_{I}-1$ real unknown [4], and solving such a system requires the inversion of a $(2N_{I}-1)\times(2N_{I}-1)$ matrix, the complexity of its solution scales with $\mathcal{O}(N_{I}^{3})$.
@@ V-A Scaling Law with Mutual Coupling
Third, to compute $\mathbb{E}[|\mathbf{z}_{RI}\Re\{\mathbf{Z}_{II}\}^{-1}\mathbf{z}_{IT}|]$ in (55), we use the eigenvalue decomposition of $\Re\{\mathbf{Z}_{II}\}^{-1}$. Since $\Re\{\mathbf{Z}_{II}\}^{-1}$ is a real symmetric matrix, we have $\Re\{\mathbf{Z}_{II}\}^{-1}=\mathbf{Q}\mathbf{\Lambda}\mathbf{Q}^{T}$, where $\mathbf{Q}\in\mathbb{R}^{N_{I}\times N_{I}}$ is an orthogonal matrix whose columns are the eigenvectors of $\Re\{\mathbf{Z}_{II}\}^{-1}$, and $\mathbf{\Lambda}=\text{diag}(\lambda_{1},\ldots,\lambda_{N_{I}})$, with $\lambda_{n_{I}}\in\mathbb{R}$ being the $n_{I}$th eigenvalue of $\Re\{\mathbf{Z}_{II}\}^{-1}$. Consequently, the random variable $R=\mathbf{z}_{RI}\Re\{\mathbf{Z}_{II}\}^{-1}\mathbf{z}_{IT}$ can be rewritten as $R=\mathbf{q}_{RI}\mathbf{\Lambda}\mathbf{q}_{IT}$, where $\mathbf{q}_{RI}=\mathbf{z}_{RI}\mathbf{Q}$ is distributed as $\mathbf{q}_{RI}\sim\mathcal{CN}(\mathbf{0},\rho_{RI}\mathbf{I})$ and $\mathbf{q}_{IT}=\mathbf{Q}^{T}\mathbf{z}_{IT}$ is distributed as $\mathbf{q}_{IT}\sim\mathcal{CN}(\mathbf{0},\rho_{IT}\mathbf{I})$ due to the orthogonality of $\mathbf{Q}$. Since $R$ can be expressed as $\mathbf{q}_{RI}\mathbf{\Lambda}\mathbf{q}_{IT}, RR$ is the sum of $N_{I}$ independent random variables, i.e., $R=\sum_{n_{I}=1}^{N_{I}}R_{n_{I}}$, where $R_{n_{I}}=[\mathbf{q}_{RI}]_{n_{I}}\lambda_{n_{I}}[\mathbf{q}_{IT}]_{n_{I}}$ has variance $\rho_{RI}\rho_{IT}\lambda_{n_{I}}^{2}$, for $n_{I}=1,\ldots,N_{I}$. Thus, according to the Lyapunov central limit theorem (CLT), $R$ is distributed as $R\sim\mathcal{CN}(0,\rho_{RI}\rho_{IT}\sum_{n_{I}=1}^{N_{I}}\lambda_{n_{I}}^{2})$. Since $\lambda_{n_{I}}^{2}$, for $n_{I}=1,\ldots,N_{I}$, are the eigenvalues of $\Re\{\mathbf{Z}_{II}\}^{-2}$, we have $\sum_{n_{I}=1}^{N_{I}}\lambda_{n_{I}}^{2}=\text{Tr}(\Re\{\mathbf{Z}_{II}\}^{-2})$, giving $R\sim\mathcal{CN}(0,\rho_{RI}\rho_{IT}\text{Tr}\left(\Re\{\mathbf{Z}_{II}\}^{-2}\right))$ and